# AnyBimanual: Transferring Unimanual Policy for General Bimanual Manipulation

### v1

@@ 4 Experiments
In this section, we first introduce the experimental setups (Section 4.1). Then, we transfer various unimanual methods to bimanual manipulation via AnyBimanual, and compare them with the state-of-the-art approaches to show the superiority (Section 4.2). We conduct an ablation study to evaluate the validity of the proposed components (Section 4.3). We further interpret the learned skill representation and decomposed volumetric representation by visualization (Section 4.4). Finally, we report the real-robot results to demonstrate the effectiveness of AnyBimanual in real-world applications (Section 4.5).
@@ 4.1 Experiment Setup
Simulation. For benchmarking, our simulation experiments are conducted on RLBench2 [31], a bimanual version extended from the widely-used RLBench [36] benchmark in prior works [35, 69, 47, 28, 63, 39]. Following the setup in [31], we utilize $12$ language-conditioned bimanual manipulation tasks varying from different challenge levels. The diverse task suite requires the agent to acquire and correctly schedule shareable skills to achieve high success rates, rather than merely imitating limited expert demonstrations. For observation, we employ six cameras with a resolution of $256\times 256$ to cover the entire workspace. During the training phase, we provide $20$ or $100$ demonstrations for each task, and we evaluate $25$ episodes per task in the testing set to mitigate the randomness.
@@ 4.2 Comparison with the State-of-the-Art Methods
In this section, we compare our AnyBimanual with previous stat-of-the-art approaches on RLBench tasksuite. Table 1 presents a comparison of the average success rates for each task and the average performance is shown in Figure 1. Our method achieves the highest overall performance, with an average success rate of $21.67$%, setting a new state-of-the-art in general bimanual manipulation. AnyBimanual leverages the knowledge distilled from unimanual models and successfully transfers it to guide general bimanual manipulation. As a result, our method outperforms the cutting-edge bimanual method PerAct2 by a significant improvement of $12.67$% on average. With a limited set of $20$ demonstrations, our method still defeats the baseline with a sizable margin of $8.00$%. Note that we aim to train multi-task agents to handle a wide range of tasks within a single model, hence the performance is lower than the reported in the original paper. Especially, we observe that AnyBimanual shows greater improvement in long-horizon tasks like put in fridge, multi-variations tasks like press buttons, and tasks demand synchronous coordinations like pick plate. Moreover, AnyBimanual enables plug-and-play transferring of PerAct-LF and RVT, which also leads to relative boosts of $25.87$% ($7.00$% to $9.00$%) and $116.65$% ($4.00$% to $8.67$%) on average. However, incorporating AnyBimanual slightly affects the performance on the simple, short-horizon Lift ball, due to the additional complexity in input processing.
@@ 4.3 Ablation Study
Skill Manager. By employing the skill manager in the experiment, we observe that the average success rate increases by $4.67$% (Row $4$ vs. Row $2$). Notably, in tasks requiring long-term manipulation within the Long category, it significantly outperforms the vanilla version, demonstrating the skill manager’s effectiveness in long-horizon tasks.
Visual Aligner. Additionally, we incorporate the spatial soft masking from the visual aligner, resulting in a substantial performance improvement of 4.34% (Row $3$ vs. Row $2$). Although the inclusion of spatial soft masking mechanism may slightly affect performance in simpler tasks due to increased input processing complexity, it leads to significant gains in overall task success. The gains are even more notable in Generalized and Sync tasks, which demand a high degree of adaptability to scene variations and synchronization for cooperative actions. The results demonstrate the visual aligner preserves the generalizability stored in the unimanual policy perfectly, facilitating seamless coordination between the arms. By integrating all techniques in our AnyBimanual, the success rate improves from $9.00$% to $21.67$% (Row $5$), validating the importance of properly mining unimanual knowledge for bimanual manipulation.
@@ 5 Conclusion
In this paper, we have introduced AnyBimanual, a framework designed to transfer pre-trained unimanual manipulation policies to multi-task bimanual manipulation with few bimanual demonstrations. We develop a skill manager to dynamically schedule skill primitives discovered from unimanual policies, enabling their effective adaptation for bimanual tasks. To address the observation discrepancies between unimanual and bimanual systems, we propose a visual aligner that generates spatial soft masks, aligning the visual embeddings of each arm with those used during the pretraining stage of the unimanual policy model. Extensive experiments across $12$ simulated and $9$ real-world tasks demonstrate the effectiveness of AnyBimanual. The limitations are discussed in the supplementary file.
@@ 7.2 Architectures
Skill Manager. We fuse voxel embeddings and language embeddings by projecting both to a hidden size of $256$ and concatenating them. After layer normalization, the combined sequence is processed by a frozen GPT-2 Transformer configured with a hidden size of $128$ and $4$ attention heads. The transformer’s output is passed through a linear layer to produce logits over 18 classes, and softmaxed to obtain probabilities. These probabilities are then used to reconstruct the skill representation via a weighted sum over the skill primitives.
@@ 8.1 Skill Clustering
The goal of the skill manager is to obtain the language embedding via linear skill representations, so that the pre-trained unimanual policy model (i.e., PerAct) can be prompted to generate feasible manipulation actions. In Figure 12, we cluster the skill representations while evaluating the last checkpoint of PerAct+AnyBimanual in RLBench2. We can observe that the learned skill representations are reasonably clustered into $18$ categories, which corresponds to the number of pre-trained unimanual tasks. We also visualize the related unimanual skill of each cluster by finding the nearest unimanual task embedding. The clustering results further interpret that the proposed skill manager enables compact skill representation learning with rich semantics for dynamic scheduling.

### v2

@@ 4 Experiments
In this section, we first introduce the experimental setups (Section 4.1). Then, we transfer various unimanual methods to bimanual manipulation via AnyBimanual, and compare them with the state-of-the-art to show uperiority (Section 4.2). We conduct an ablation study to evaluate validity of the proposed components (Section 4.3). We further interpret learned skill representation and decomposed volumetric representation by visualization (Section 4.4). Finally, we report real-robot results to demonstrate effectiveness of AnyBimanual in real-world applications (Section 4.5).
@@ 4.1 Experiment Setup
Simulation. For benchmarking, our simulation experiments are conducted on RLBench2 [31], a bimanual version extended from the widely-used RLBench [36] benchmark in prior works [35, 69, 47, 28, 63, 39]. Following the setup in [31], we utilize $12$ language-conditioned bimanual manipulation tasks varying from different challenge levels. The diverse task suite requires the agent to acquire and correctly schedule shareable skills to achieve high success rates, rather than merely imitating limited expert demonstrations. For observation, we employ six cameras with a resolution of $256\times 256$ to cover the entire workspace. During the training phase, we provide $20$ or $100$ demonstrations for each task, and we evaluate $100$ episodes per task in the testing set to mitigate the randomness.
@@ 4.2 Comparison with the State-of-the-Art Methods
In this section, we compare our AnyBimanual with previous stat-of-the-art approaches on RLBench tasksuite. Table 1 presents a comparison of the average success rates for each task and the average performance is shown in Figure 1. Our method achieves the highest overall performance, with an average success rate of $32.00$%, setting a new state-of-the-art in general bimanual manipulation. AnyBimanual leverages the knowledge distilled from unimanual models and successfully transfers it to guide general bimanual manipulation. As a result, our method outperforms the cutting-edge bimanual method PerAct2 by a significant improvement of $17.33$% on average. With a limited set of $20$ demonstrations, our method still defeats the baseline with a sizable margin of $10.50$%. Note that we aim to train multi-task agents to handle a wide range of tasks within a single model, hence the performance is lower than the reported in the original paper. Especially, we observe that AnyBimanual shows greater improvement in long-horizon tasks like put in fridge, multi-variations tasks like press buttons, and tasks demand synchronous coordinations like straighten rope. Moreover, AnyBimanual enables plug-and-play transferring of PerAct-LF and RVT, which also leads to relative boosts of $72.76$% ($4.92$% to $8.50$%) and $39.41$% ($10.58$% to $14.75$%) on average. However, incorporating AnyBimanual slightly affects the performance on the simple, short-horizon Lift ball, due to the additional complexity in input processing.
@@ 4.3 Ablation Study
Skill Manager. By employing the skill manager in the experiment, we observe that the average success rate increases by $8.92$% (Row $4$ vs. Row $2$). Notably, in tasks requiring long-term manipulation within the Long category, it significantly outperforms the vanilla version, demonstrating the skill manager’s effectiveness in long-horizon tasks.
Visual Aligner. Additionally, we incorporate the spatial soft masking from the visual aligner, resulting in a substantial performance improvement of 3.00% (Row $3$ vs. Row $2$). Although the inclusion of spatial soft masking mechanism may slightly affect performance in simpler tasks due to increased input processing complexity, it leads to significant gains in overall task success. The gains are even more notable in Generalized and Sync tasks, which demand a high degree of adaptability to scene variations and synchronization for cooperative actions. The results demonstrate the visual aligner preserves the generalizability stored in the unimanual policy perfectly, facilitating seamless coordination between the arms. By integrating all techniques in our AnyBimanual, the success rate improves from $14.67$% to $32.00$% (Row $5$), validating the importance of properly mining unimanual knowledge for bimanual manipulation.
@@ 5 Conclusion
In this paper, we have introduced AnyBimanual, a framework designed to transfer pretrained unimanual manipulation policies to multi-task bimanual manipulation with few bimanual demonstrations. We develop a skill manager to dynamically schedule skill primitives discovered from unimanual policies, enabling their effective adaptation for bimanual tasks. To address the observation discrepancies between unimanual and bimanual systems, we propose a visual aligner that generates spatial soft masks, aligning the visual embeddings of each arm with those used during the pretraining stage of the unimanual policy model. Extensive experiments across $12$ simulated and $9$ real-world tasks demonstrate the effectiveness of AnyBimanual. The limitations are discussed in the supplementary file.
@@ 7.2 Architectures
Skill Manager. We fuse voxel embeddings and language embeddings by projecting both to a hidden size of $256$ and concatenating them. After layer normalization, the combined sequence is processed by a frozen GPT-2 Transformer configured with a hidden size of $256$ and $8$ attention heads. The transformer’s output is passed through a linear layer to produce logits over 18 classes, and softmaxed to obtain probabilities. These probabilities are then used to reconstruct the skill representation via a weighted sum over the skill primitives.
@@
7.3 Skill Primitives
The skill primitives use CLIP embeddings of unimanual language templates (e.g., ‘open the   drawer’) from Table 1 of the supplementary file, as we expect them to represent foundational motions that generalize across object categories in unseen scenarios.
@@ 8.1 Skill Clustering
The goal of the skill manager is to obtain the language embedding via linear skill representations, so that the pre-trained unimanual policy model (i.e., PerAct) can be prompted to generate feasible manipulation actions. In Figure 12, we cluster the skill representations while evaluating the last checkpoint of PerAct+AnyBimanual in RLBench2. We can observe that the learned skill representations are reasonably clustered into $18$ categories, which corresponds to the number of pretrained unimanual tasks. We also visualize the related unimanual skill of each cluster by finding the nearest unimanual task embedding. The clustering results further interpret that the proposed skill manager enables compact skill representation learning with rich semantics for dynamic scheduling.
@@ 8.2 Rollout Comparisons
8.3 Visualization
Skill Manager. We visualize the skill manager in two challenging tasks requiring heavy coordination. (Figure 15) AnyBimanual completes these tasks by scheduling skill representations across arms simultaneously and dynamically.
Figure 15: Visualization of Skill Manager.
Viusal Aligner. By predicting spatial soft masks for voxel observation, the visual aligner encourages both arms to focus on the interaction and ignore irrelevant parts. (Figure 16)
@@ 8.4 Error Modes
Figure 16: Visualization of Visual Aligner.
8.5 Bimanual Decomposability
We complement a heuristic measurement to assess the feasibility of decomposing a given bimanual task with a learned skill manager at the semantic level. Specifically, higher entropy in the predicted combination weights indicates greater difficulty in decomposition with certain unimanual primitives. In Figure 17, this metric strongly correlates with final success rates in four simulation tasks, delineating the boundary.
Figure 17: Bimanual Decomposability.